Given the total cost function; Q = 0.1Q3- 3Q2+ 60Q + 5400

A. Calculate each average cost function and show it graphically

B. What is the level output at the minimum level of average variable cost?

c. What is the level of output a1 the minimum point of average total cost and

what is the trend or average variable cost at this point?

D. What is the fixed cost and what does it mean?


1
Expert's answer
2021-09-21T15:04:39-0400

A. ATC=TC/Q=0.1Q2-3Q+60+5400/Q AFC=5400/Q; AVC=0.1Q2-3Q+60

B. Average variable cost reaches the minimum at the point where AVC'=0. AVC'=0.2Q-3=0; Q=15.

C. Similarly ATC'=0; ATC'=0.2Q-3-5400/Q^2=0; 0.2Q3-3Q2-5400=0; Q=36. Average variable cost increases at this point

D. The fixed cost is 5400; this cost exists even when the firm doesn't produce anything.